OpenStudy (anonymous):

(a^x + 1)/( b^x-1) = 10; what is x?

OpenStudy (unklerhaukus):

what are \(a\) and \(b\)

OpenStudy (anonymous):

There are three variables how one can find three with just one equation??

OpenStudy (anonymous):

i think a and b are constants if so then hard to get x

OpenStudy (unklerhaukus):

are they integers

OpenStudy (anonymous):

@radithya

OpenStudy (anonymous):

Somehow I got: \(a = 2, b = \sqrt{3}, x = 2 \)

OpenStudy (unklerhaukus):

i thought the solution might be \(x=x(a,b)\) how could you solve three variables from on equation @waterineyes ?

OpenStudy (anonymous):

= (x+1) log a - (x-1) log b = log 10

OpenStudy (anonymous):

i think its \[\huge \frac{a^x+1}{b^x-1}=10\]

OpenStudy (jiteshmeghwal9):

\[\frac{a^x+1}{b^x-1}=\frac{10}{1}\]\[(a^x+1)=10(b^x-1)\[a^x+1=10b^x-1\]

OpenStudy (unklerhaukus):

@waterineyes 5/2≠10

OpenStudy (anonymous):

Sorry 20/2 is to be taken...

OpenStudy (unklerhaukus):

\[(a^x+1)=10(b^x−1)\]\[a^x+1=10b^x-10\]\[a^x=10b^x-11\]

OpenStudy (jiteshmeghwal9):

ya! thats wht i wants to say

OpenStudy (anonymous):

\[\frac{a^x + 1}{b^x - 1} = \frac{50}{5}\implies a^x = 49 \implies a = 7 \quad and \quad x = 2\] So, \(b = \sqrt{6}\) \(a= 7, b = \sqrt{6}, x = 2\) And It is also true that solutions can be many....

OpenStudy (unklerhaukus):

\[x=\log_a\left(10b^x-11\right)\]

OpenStudy (anonymous):

There is x in the right hand side too.. It is not right...

OpenStudy (anonymous):

For calculating x, it will be like: x = something And this something must not contain x...

OpenStudy (anonymous):

You are going wrong @jiteshmeghwal9 You have calculated one variable from one equation and are going to substitute that variable in the same equation...

OpenStudy (anonymous):

guys i think its impossible to calculate x in the form x=x(a,b) !!!!!! :D

OpenStudy (anonymous):

Now you will get 0 @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

ohhh ! yeah

OpenStudy (anonymous):

im sure****

OpenStudy (anonymous):

There will come \(10b^x - 10\) in the numerator @jiteshmeghwal9

OpenStudy (unklerhaukus):

infinite lines of solutions if a,b<7

OpenStudy (anonymous):

thanks, its ok even if i don't get the answer